How much charge flows from a 12.0 V battery when it is connected to a 9.00 microfarad capacitor?


1
Expert's answer
2022-03-11T08:28:37-0500

The charge is given as the product of capacitance and voltage:


"Q= CV = 9\times 10^{-6}F\cdot 12V = 108\times 10^{-6}C = 1.08\times 10^{-4}C"

Answer. "1.08\times 10^{-4}C".